# The need for accuracy and smoothness in numerical simulations

### v1

@@ Abstract
We consider the problem of estimating the error when solving a system of differential algebraic equations. Richardson extrapolation is a classical technique that can be used to judge when computational errors are irrelevant and estimate the discretization error. We have simulated molecular dynamics with constraints using the GROMACS library and found that the output is not always amenable to Richardson extrapolation. We derive and illustrate Richardson extrapolation using a variety of numerical experiments. We identify two necessary conditions that are not always satisfied by the GROMACS library.
@@ 1 Motivation
Consider the problem simulating the motion of a system of atoms moving in a force field subject to a set of constraints. In this case, Newton’s 2nd law takes the form of the following system of differential algebraic equations
@@ 2 Theory
We shall now show how to demonstrate the existence of an asymptotic error expansion experimentally. We define Richardson’s fraction $F_{h}$ using the expression
@@
$F_{h}=\frac{A_{2h}-A_{4h}}{A_{h}-A_{2h}}=2^{p}\left[\frac{1+2^{m}\nu h^{m}+\phi(h)}{1+\nu h^{m}+\phi(2h)}\right].$ (26)
@@
$F_{h}=2^{p}\left(1+\frac{(2^{m}-1)\nu h^{m}}{1+g(h)}+\frac{\phi(h)-\phi(2h)}{1+g(h)}\right)\rightarrow 2^{p},\quad h\rightarrow 0_{+}$ (29)
@@
We conclude that if $T-A_{h}$ satisfies the asymptotic error expansion (10), then the order of the primary error term can be determined as the limit
$p=\underset{h\rightarrow 0_{+}}{\lim}F_{h}$ (32)
@@
is a good approximation for $h$ sufficiently small. In particular, we note that the high-hand side of equation 33 is a linear function of $\log(h)$ with slope $m$.
@@ 3 Elementary examples
The theory applies to the difference $T-A_{h}$ between the target value $T$ and the exact value of the approximation $A_{h}$. In practice, the computed value $\hat{A}_{h}$ is different from the exact value $A_{h}$. However, it is often possible to assert that the computational error $A_{h}-\hat{A}_{h}$ is irrelevant and issue precise estimates of the error $T-\hat{A}_{h}$. We shall now demonstrate the procedure. To this end, we consider the familiar problem of computing definite integrals
@@ 4.1 A successful application of the theory
The script maxrange_rk1 models a shell as a point particle moving in a plane subject to Earth’s standard gravity and the international standard atmospheric model. Each trajectory is integrated using Euler’s explicit method (’rk1’) and all but the final step has the same size $h$. The final step is adjusted to place the shell on the ground. The drag functions are interpolated from tables using cubic spline interpolation. The function plot_shells will plot the drag coefficients for the different shells as a function of the Mach number. For each drag coefficient our target value $T$ is the maximum range of the shell as the elevation of the howitzer varies continuously from 00 to $\frac{\pi}{2}$. For each drag coefficient, we compute $12$ different approximations $A_{h_{k}}$ of $T$ using the step size $h_{k}=2^{3-k}$ seconds, where $k\in\{1,2,\dots,12\}$. For each drag coefficient and for each value of the time step $h$, a range function is defined which returns the range of the shell as a function of the howitzers elevation $\theta$. The range functions are unimodal and the maximum range is found using the golden section search algorithm. The initial search bracket is $[0,\pi/2$] and this bracket is systematically reduced in length until it is shorter than $\frac{\pi}{2}u$. The script will either read the raw data from a file or generate it from scratch. In any case, the script produces several figures and tables including Figure 1 and Figure 3.
@@
These two figures represent calculations where the final time step is computed with an error that is bounded by $uh_{k}$ where $u=2^{-53}$ is the double precision unit roundoff. Figure 1 lists the maximum range and the corresponding error estimate for each of the 6 shell types in our library using a time step of $h=2^{-9}s$. In each case the error estimate suggests that the computed range is exact to the number of figures shown. In particular, we see that a G7 type shell achieves a maximum range of 17.5 km and all other shells have ranges that are less than 16.9 km. However, it is a fallacy to conclude anything on the basis of this table alone. In each case, we need to assert that we are inside the asymptotic range and that the error estimates are reliable. To this end, we examine the evolution of Richardson’s fraction for the maximum range of each shell, see Figure 3. For each of the 3 different drag coefficients shown we see that the evolution of Richardson’s fraction supports an asymptotic error expansion with $(p,q)=(1,2)$. This result is consistent with the use of Euler’s explicit method which is 1st order accurate in the time step [^1] The remaining figures are similar and have been omitted to save space.. We observe that for each drag coefficient, $k=12$ is still inside the asymptotic range and we have no reason to doubt the magnitude of the error estimate. We conclude that the best model for the D-20 howitzer is in fact the G7 shell.
@@ 5.1 The need for sufficient accuracy
When the tolerance is large, see Figure 6, there is no evidence that an asymptotic error expansion exists and there is no reason to trust Richardson’s error estimate. When the tolerance is small, see Figure 7, the fact that $\hat{F}_{h}$ approaches $2^{p}$ suggests that an asymptotic error expansion exists, but it is not trivial to determine an asymptotic range. Regardless, it is clear that if we do not know which tolerance is sufficient and if an error estimate is required, then our safest course of action is to solve all equations as accurately as the hardware will allow.
@@ 6 Conclusion
A central task in the field of computational science is to fit a system of DAEs to the results of a physical experiment. It is often possible to assert when the computational errors are irrelevant and estimate the discretization error using Richardson extrapolation. This classical technique hinges on the existence of an asymptotic error expansion. However, if the functions that describe our problem are not sufficiently differentiable or if the computational error is not sufficiently small, then we cannot estimate the discretization error and we lose the ability to evaluate our model. In the absence of further analysis, the best strategy is therefore to use functions that are smooth and solve all central equations as accurately as possible.
@@ 6.0.1 Acknowledgments
The authors would like to thank Jesús Alastruey-Benedé, Pablo Ibáñez and Pablo García-Risueño for stimulating discussions on the subject matter. The first author is supported by eSSENCE, a collaborative e-Science programme funded by the Swedish Research Council within the framework of the strategic research areas designated by the Swedish Government. This work has been partially supported by the Spanish Ministry of Science and Innovation MCIN/AEI/10.13039/501100011033 (grant PID2022-136454NB-C22), and by Government of Aragon (T58_23R research group).

### v2

@@ Abstract
We consider the problem of estimating the error when solving a system of differential algebraic equations. Richardson extrapolation is a classical technique that can be used to judge when computational errors are irrelevant and estimate the discretization error. We have simulated molecular dynamics with constraints using the GROMACS library and found that the output is not always amenable to Richardson extrapolation. We derive and illustrate Richardson extrapolation using a variety of numerical experiments. We identify two necessary conditions that are not always satisfied by the GROMACS library. [^1] This is a preprint of a paper accepted by the conference PPAM-2024. The paper is to appear in Springer’s LNCS series.
@@ 1 Motivation
Consider the problem of simulating the motion of a system of atoms moving in a force field subject to a set of constraints. In this case, Newton’s 2nd law takes the form of the following system of differential algebraic equations
@@ 2 Theory
We shall now demonstrate how the values of $p$ and $q$ can be determined by observing the exact values of $A_{h}$ for different values of $h$. We define Richardson’s fraction $F_{h}$ using the expression
@@
$F_{h}=\frac{A_{2h}-A_{4h}}{A_{h}-A_{2h}}=2^{p}\left[\frac{1+2^{m}\nu h^{m}+\phi(2h)}{1+\nu h^{m}+\phi(h)}\right].$ (26)
@@
$F_{h}=2^{p}\left(1+\frac{(2^{m}-1)\nu h^{m}}{1+g(h)}+\frac{\phi(2h)-\phi(h)}{1+g(h)}\right)\rightarrow 2^{p},\quad h\rightarrow 0_{+}$ (29)
@@
We conclude that if $T-A_{h}$ satisfies the asymptotic error expansion (10), then the order of the primary error term can be determined from the limit
$2^{p}=\underset{h\rightarrow 0_{+}}{\lim}F_{h}$ (32)
@@
is a good approximation for $h$ sufficiently small. In particular, we note that the right-hand side of equation 33 is a linear function of $\log(h)$ with slope $m$.
@@ 3 Elementary examples
The theory applies to the difference $T-A_{h}$ between the target value $T$ and the exact value of the approximation $A_{h}$. In practice, the computed value $\hat{A}_{h}$ is different from the exact value $A_{h}$. However, as we shall demonstrate shortly, is often possible to assert that the computational error is irrelevant and estimate the error $T-\hat{A}_{h}$ accurately. We begin by considering the familiar problem of computing definite integrals
@@ 4.1 A successful application of the theory
The script maxrange_rk1 models a shell as a point particle moving in a plane subject to Earth’s standard gravity and the international standard atmospheric model. Each trajectory is integrated using Euler’s explicit method (’rk1’) and all but the final step has the same size $h$. The final step is adjusted to place the shell on the ground. The drag functions are interpolated from tables using cubic spline interpolation. The function plot_shells will plot the drag coefficients for the different shells as a function of the Mach number. For each drag coefficient our target value $T$ is the maximum range of the shell as the elevation of the howitzer varies continuously from 00 to $\frac{\pi}{2}$. For each drag coefficient, we compute $12$ different approximations $A_{h_{k}}$ of $T$ using the step size $h_{k}=2^{3-k}$ seconds, where $k\in\{1,2,\dots,12\}$. For each drag coefficient and for each value of the time step $h$, a range function is defined which returns the range of the shell as a function of the howitzers elevation $\theta$. The range functions are unimodal and the maximum range is found using the golden section search algorithm. The initial search bracket is $[0,\pi/2$] and this bracket is systematically reduced in length until it is shorter than $\frac{\pi}{2}u$. The script will either read the raw data from a file or generate it from scratch. In any case, the script produces several figures and tables including Table 1 and Figure 3.
@@
These two figures represent calculations where the final time step is computed with an error that is bounded by $uh_{k}$ where $u=2^{-53}$ is the double precision unit roundoff. Table 1 lists the maximum range and the corresponding error estimate for each of the 6 shell types in our library using a time step of $h=2^{-9}s$. In each case the error estimate suggests that the computed range is exact to the number of figures shown. In particular, we see that a G7 type shell achieves a maximum range of 17.5 km and all other shells have ranges that are less than 16.9 km. However, it is a fallacy to conclude anything on the basis of this table alone. In each case, we need to assert that we are inside the asymptotic range and that the error estimates are reliable. To this end, we examine the evolution of Richardson’s fraction for the maximum range of each shell, see Figure 3. For each of the 3 different drag coefficients shown we see that the evolution of Richardson’s fraction supports an asymptotic error expansion with $(p,q)=(1,2)$. This result is consistent with the use of Euler’s explicit method which is 1st order accurate in the time step [^2] The remaining figures are similar and have been omitted to save space.. We observe that for each drag coefficient, $k=12$ is still inside the asymptotic range and we have no reason to doubt the magnitude of the error estimate. We conclude that the best model for the D-20 howitzer is in fact the G7 shell.
@@ 5.1 The need for sufficient accuracy
When the tolerance is $\tau=2^{-11}$, see 6, there is no evidence that an asymptotic error expansion exists and there is no reason to trust Richardson’s error estimate. When the tolerance is $\tau=2^{-25}$, see Figure 7, the fact that $\hat{F}_{h}$ approaches $2^{p}$ suggests that an asymptotic error expansion exists, but it is not trivial to determine an asymptotic range. Regardless, it is clear that if we do not know which tolerance is sufficient and if an error estimate is required, then our safest course of action is to solve all equations as accurately as the hardware will allow.
@@ 6 Conclusion
A central task in the field of computational science is to fit a system of DAEs to the results of a physical experiment. It is often possible to assert when the computational errors are irrelevant, estimate the discretization error using Richardson extrapolation. This classical technique hinges on the existence of an asymptotic error expansion. However, if the functions that describe our problem are not sufficiently differentiable or if the computational error is not sufficiently small, then we cannot estimate the discretization error and we lose the ability to evaluate our model. In the absence of further analysis, the best strategy is therefore to use functions that are smooth and solve all central equations as accurately as possible.
@@ 6.0.1 Acknowledgments
We would like to thank the reviewers for their work which allowed us to improve the manuscript. We would like to thank Jesús Alastruey-Benedé, Pablo Ibáñez and Pablo García-Risueño for stimulating discussions on the subject matter. The first author is supported by eSSENCE, a collaborative e-Science programme funded by the Swedish Research Council within the framework of the strategic research areas designated by the Swedish Government. This work has been partially supported by the Spanish Ministry of Science and Innovation MCIN/AEI/10.13039/501100011033 (grant PID2022-136454NB-C22), and by Government of Aragon (T58_23R research group).